3.6gram of a mixture of nacl and kcl give 7.74gram of agcl with agno3 find out the percentage of each salt in the mixture. Plzzz help me.. I also got answers... .. but I dont know how. To solve this... Answers are.. . %nacl=42.5 and %kcl=57.5

Grade:11

1 Answers

varun
6 years ago
HELLO SHUBH,
do not worry i will help with your problem
its little tricky but simple in solving
i will explain you as follows please look out
since moles of AgCl = 7.74/143.5 
moles of Cl- will aslo be 7.74/143.5
NOW,
let no of moles of KCl salt be = a moles
let no of moles of NaCl salt be = b moles
so,
a+b = 7.74/143.5
        = 0.05 M
next,
3.6 = 74.5a + 58.5b
so after solving the above equation,
a=0.028
b = 0.026
next,
mass of KCl = 78.5 X a
                    = 78.5 x 0.028
                     =2.06gms
mass of NaCl = 58.5 X b
                     = 58.5 x 0.026
                        = 1.53gms\
therefore,
percentage of NaCl = 1.53/3.6 X 100
                            =42.5%
percentage of KCl = 2.06/3.6 X 100
                            = 57.4%
approve me if my solution is right
THANK YOU